Find the power set of each of these sets, where a and b are distinct elements. a) {a} b) {a, b} c) {∅,{∅}}


1
Expert's answer
2022-03-29T17:32:02-0400

"Power\\,\\,set\\,\\,is\\,\\,the\\,\\,set\\,\\,of\\,\\,all\\,\\,subsets\\\\a:2^{\\left\\{ a \\right\\}}=\\left\\{ \\emptyset ,\\left\\{ a \\right\\} \\right\\} \\\\b:2^{\\left\\{ a,b \\right\\}}=\\left\\{ \\emptyset ,\\left\\{ a \\right\\} ,\\left\\{ b \\right\\} ,\\left\\{ a,b \\right\\} \\right\\} \\\\c:2^{\\left\\{ \\emptyset ,\\left\\{ \\emptyset \\right\\} \\right\\}}=\\left[ \\begin{array}{c}\tin\\,\\,the\\,\\,previous\\,\\,answer\\\\\tsubstitute\\,\\,a\\,\\,with\\,\\,\\emptyset\\\\\tand\\,\\,b\\,\\,with\\,\\,\\left\\{ \\emptyset \\right\\}\\\\\\end{array} \\right] =\\left\\{ \\emptyset ,\\left\\{ \\emptyset \\right\\} ,\\left\\{ \\left\\{ \\emptyset \\right\\} \\right\\} ,\\left\\{ \\emptyset ,\\left\\{ \\emptyset \\right\\} \\right\\} \\right\\}"